OpenStudy (anonymous):

1/4 plus 1/4 plus 3/5 plus 3/5 ??

OpenStudy (anonymous):

17/10

OpenStudy (anonymous):

Pay attention....

OpenStudy (aroub):

1/4+1/4 +3/5+3/5 2/4+6/5 2/4*5 +6/5*4 10/20 +24/20 34/20 17/10

OpenStudy (anonymous):

im trying to figure this stuff out but it takes a while for me to get it. i have ADD. and ive always been slow at math but all the other subjects im good at

OpenStudy (anonymous):

If u have add, u have subtract...

OpenStudy (anonymous):

do you know what im talkin bout? i mean attention defaset disorder or how ever you spell it

OpenStudy (anonymous):

But u only have this disorder when doing math,lol:-)

OpenStudy (anonymous):

no its just that with other subjects i catch on a lil faster but math always takes a while for me

OpenStudy (anonymous):

Subtraction is negative addition.

OpenStudy (anonymous):

i think ill sign up for a tutor or somethin..

OpenStudy (anonymous):

No, no just practice, lots of (free9 tutors here...

OpenStudy (anonymous):

yea true

OpenStudy (anonymous):

Next one u post, put your own try, never mind if it is wrong...

OpenStudy (anonymous):

also pink sweety is posting similar, u can look at them as well....

OpenStudy (anonymous):

Maybe u can do hers...

OpenStudy (anonymous):

ok this one sais monica spent 3/4 hours listing to to tapes of beethoven and brahms. she spent 1/5hours listing to beethoven. how many hours were spent listing to brahms

OpenStudy (anonymous):

Go on...

OpenStudy (anonymous):

so i do 1/5 - 3/4 right or no?

OpenStudy (anonymous):

beethoven and brahms - beethoven

OpenStudy (anonymous):

do i subtract the two fractions?

OpenStudy (anonymous):

19/20

OpenStudy (anonymous):

beethoven and brahms - beethoven = ?

OpenStudy (anonymous):

i guess its the musican beethoven

OpenStudy (anonymous):

OK, now swap the names for the numbers...

OpenStudy (anonymous):

3/4 divided by 1/5?? or no

OpenStudy (anonymous):

- (minus)

OpenStudy (anonymous):

and i have to find the lcd first right?

OpenStudy (anonymous):

Step at a time....

OpenStudy (anonymous):

So brahms and beethoven =?

OpenStudy (anonymous):

cause i know the lcd has to be 20

OpenStudy (anonymous):

would the answer be 1/10?

OpenStudy (anonymous):

3/4 = ?/20

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

cause 5 and 4 is 20

OpenStudy (anonymous):

3*5/4*5 = 15/20

OpenStudy (anonymous):

Now 1/5 = ?/20

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Correct. So u have 15/20 and 4/20 is ?

OpenStudy (anonymous):

Sorry, minus not plus....

OpenStudy (anonymous):

11/20?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

So u see what the LCD does, it is converting the bits so that they have the same number on the bottom

OpenStudy (anonymous):

yea im startin to understand more its just takin a while

OpenStudy (anonymous):

After u did many, it becomes much easier..

OpenStudy (anonymous):

i apprciate the help

OpenStudy (anonymous):

Your welcome.